OpenStudy (anonymous):

A bat hits a moving baseball. If the bat delivers a net eastward impulse of 0.5 N-s and the ball starts with an initial horizontal velocity of 3.8 m/s to the west and leaves with a 4.6 m/s velocity to the east, what is the mass of the ball (in grams)?

OpenStudy (anonymous):

I think its 0.6

OpenStudy (anonymous):

since p=m(vf-vi)

OpenStudy (anonymous):

\[F \Delta t = m(v _{f} - v _{i})\]solve for m by plugging the numbers in

OpenStudy (anonymous):

isn't p 0.5?

OpenStudy (anonymous):

\[F \Delta t \]is the impulse

OpenStudy (anonymous):

since p is impulse?

OpenStudy (anonymous):

no, Delta p is the impulse = F delta t

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

So Fdeltat is 0.5N-s?

OpenStudy (anonymous):

So 0.5N-s=m(4.6-(-3.8)) then solve for m which is 0.6

OpenStudy (anonymous):

yes, and that = m (delta v)... solve for m

OpenStudy (anonymous):

k just making sure. :) thank you!

OpenStudy (anonymous):

I think the ans is 0.625 kg